# Copula‐Stein Discrepancy: A Generator‐Based Stein Operator for Archimedean Dependence

### v1

@@ Abstract
Kernel Stein discrepancies (KSDs) have become a principal tool for goodness-of-fit testing, but standard KSDs are often insensitive to higher-order dependency structures, such as tail dependence, which are critical in many scientific and financial domains. We address this gap by introducing the Copula-Stein Discrepancy (CSD), a novel class of discrepancies tailored to the geometry of statistical dependence. By defining a Stein operator directly on the copula density, CSD leverages the generative structure of dependence, rather than relying on the joint density’s score function. For the broad class of Archimedean copulas, this approach yields a closed-form Stein kernel derived from the scalar generator function. We provide a comprehensive theoretical analysis, proving that CSD (i) metrizes weak convergence of copula distributions, ensuring it detects any mismatch in dependence; (ii) has an empirical estimator that converges at the minimax optimal rate of $O_{P}(n^{-1/2})$; and (iii) is provably sensitive to differences in tail dependence coefficients. The framework is extended to general non-Archimedean copulas, including elliptical and vine copulas. Computationally, the exact CSD kernel evaluation scales linearly in dimension, while a novel random feature approximation reduces the $n$-dependence from quadratic $O(n^{2})$ to near-linear $\tilde{O}(n)$, making CSD a practical and theoretically principled tool for dependence-aware inference.
1.  Introduction
@@
This work positions CSD as a principled, practical, and powerful tool for dependence-aware model diagnostics and goodness-of-fit testing. The remainder of the paper is structured as follows. Section 2 discusses related work. Section 3 provides background on KSDs and copulas. Section 4 introduces all the notations and assumptions used throughout the paper. Section 5 formally introduces the CSD, its operator, and kernel. Section 6 establishes its key theoretical properties. Section 7 extends the framework to general copulas, and Section 8 details its computational properties and scalable algorithms.
2.  Related Work
@@ 2.2 Goodness-of-Fit Testing for Copulas
CSD contributes a new class of GoF tests to this literature with several key advantages. First, as a discrepancy metric, its value provides a quantitative measure of misspecification, whereas many classical tests only yield a p-value. Second, our framework avoids the need for smoothing parameters or potentially invalid parametric bootstrap procedures (Genest and Rémillard, 2008) by leveraging the wild bootstrap techniques common in the KSD literature. Most importantly, CSD is specifically designed to be sensitive to features like tail dependence, a known challenge highlighted in financial applications (Malevergne and Sornette, 2003). While specialized tests for extreme-value dependence exist (Kojadinovic and Yan, 2011; Genest and Segers, 2009), CSD provides a general framework with theoretical guarantees on its sensitivity to tail coefficients for any copula family.
@@ 2.3 Copula Modeling and Diagnostics
3.  Background: Copulas and Stein’s Method
@@ 3.2 Kernel Stein Discrepancy
4.  Standing Assumptions and Notation
What this section does.
We state the standing assumptions (the regularity conditions used throughout the paper) and fix the notation used in all sections. The results that follow cite these as “Assumptions A1–A5” (and A2 when needed).
Distributions and Samples.
A probability distribution is $P$ with density $p$. The target (null) copula distribution is $C_{0}$ with density $c_{0}$; a generic copula is $C$ with density $c$. The dd–dimensional unit hypercube is $0,1]^{d}$. An i.i.d. sample of size $n$ from a copula is $\{U_{i}\}_{i=1}^{n}$, $U_{i}\in[0,1]^{d}$. The empirical measure is $C_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{U_{i}}$. Expectations are $\mathbb{E}[\cdot]$.
Copula Framework.
A copula is $C(u)$ with $u=(u_{1},\dots,u_{d})\in[0,1]^{d}$. For Archimedean copulas, $\varphi:(0,1]\to[0,\infty)$ is the generator, $\psi=\varphi^{-1}$, and $\varphi^{(k)}$ denotes the $k$th derivative. Parametric dependence parameters are denoted $\theta$. The copula score is $s(u)=\nabla\log c(u)$. Lower/upper tail coefficients are $\lambda_{L},\lambda_{U}$. For Gaussian (elliptical) copulas, $\Sigma$ is the correlation matrix and $\Phi^{-1}$ the standard normal quantile.
RKHS/Stein Framework.
The base kernel is $k(u,v)$ on $0,1]^{d}\times[0,1]^{d}$ with scalar RKHS $\mathcal{H}$ and vector RKHS $\mathcal{H}^{d}$. A vector test function is $g(u)\in\mathcal{H}^{d}$. The Copula-Stein operator at target $C_{0}$ is
$\mathcal{T}_{C_{0}}g(u)=\sum_{j=1}^{d}\!\Big[\partial_{u_{j}}g_{j}(u)+g_{j}(u)\,\partial_{u_{j}}\log c_{0}(u)\Big].$
@@ 4.  Standing Assumptions and Notation
$\widehat{\mathrm{CSD}}_{n}^{2}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}).$
Unless noted, each component $g_{j}$ vanishes on the faces $\{u_{j}=0\}$ and $\{u_{j}=1\}$.
Definition 4.1 (Copula Stein discrepancy)
@@
$\mathrm{CSD}(Q,C_{0})\;\equiv\;\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\,\mathbb{E}_{U\sim Q}\big[\,\mathcal{T}_{C_{0}}g(U)\,\big].$
Under Assumption 1 (A1–A4), we have $\mathrm{CSD}(C_{0},C_{0})=0$ by Lemma 1.
@@
Assumption 1 (Regularity Conditions)
@@ 3. A3:
    (Kernel Properties) $k$ is continuous, characteristic, and $C^{2}$ on $(0,1)^{d}\times(0,1)^{d}$, with $k$ and its partials up to order 2 uniformly bounded on compact subsets and the corresponding RKHS is closed under first-order partial derivatives on $(0,1)^{d}$.
@@ 4. A4:
    (Boundary Vanishing) $\mathcal{H}^{d}\subset C^{1}_{0}((0,1)^{d};\mathbf{R}^{d})$, i.e., $g\in\mathcal{H}^{d}$ vanishes at the boundary so integration-by-parts has no boundary term.
@@ 5. A5:
Strengthening used only for metrization.
When needed for metrization, we additionally assume:
Assumption 2 (Boundary rates for metrization)
@@
Estimator (fixed):
We use the V-statistic form throughout:
$\widehat{\mathrm{CSD}}_{n}^{2}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}),\quad U_{i}\in[0,1]^{d}.$
Archimedean score sum:
Let $\varphi:[0,1]\to\mathbb{R}$ be the generator/score map; we write
$t(u)\;=\;\sum_{k=1}^{d}\varphi(u_{k})\,.$
Random features (Section 8):
We denote by $m\in\mathbb{N}$ the number of random features and by $\phi:\,[0,1]^{d}\to\mathbb{R}^{m}$ the random feature map used to approximate $k_{C_{0}}$. Given a draw $\omega\sim\Pi$ (feature distribution), write $\phi(u)=\phi(u;\omega_{1:m})$ for the $m$-dimensional feature vector.
We use a random-feature witness defined in Section 8.
5.  The Copula-Stein Discrepancy
The central goal of this section is to construct a dependence-aware goodness-of-fit metric relative to a target copula $C_{0}$. Throughout, we work under Assumptions 1 (and, when invoked for metrization results, Assumption 2).
@@ 5.1 From IPM to a Computable Kernel
Following the framework of integral probability metrics (IPMs) (Müller, 1997; Zolotarev, 1983), we define a discrepancy as the worst-case difference in expectation over a class of test functions $g\in\mathcal{H}^{d}$:
$\text{CSD}(C,C_{0})=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\left|\mathbb{E}_{U\sim C}[\mathcal{T}_{C_{0}}g(U)]-\mathbb{E}_{U\sim C_{0}}[\mathcal{T}_{C_{0}}g(U)]\right|.$ (2)
@@
Definition 5.1 (Copula-Stein Operator)
@@
Lemma 1 (Stein identity and boundary conditions)
@@
$\mathbb{E}_{U\sim C_{0}}\!\big[\mathcal{T}_{C_{0}}g(U)\big]=0,\quad\mathcal{T}_{C_{0}}g:=\sum_{j=1}^{d}\big(\partial_{u_{j}}g_{j}+s_{j}g_{j}\big),\ \ s_{j}=\partial_{u_{j}}\log c_{0}.$
Proof. For each $j$,
$\int\partial_{u_{j}}g_{j}(u)\,c_{0}(u)\,du=\Big[g_{j}(u)\,c_{0}(u)\Big]_{u_{j}=0}^{u_{j}=1}-\int g_{j}(u)\,\partial_{u_{j}}c_{0}(u)\,du.$
Boundary vanishing implies the bracketed term is 0, and $\partial_{u_{j}}c_{0}=s_{j}c_{0}$. Summing over $j$ yields $\int\mathcal{T}_{C_{0}}g\,c_{0}\,du=0$, i.e. $\mathbb{E}_{C_{0}}[\mathcal{T}_{C_{0}}g(U)]=0$.
Using Lemma 1, the zero-mean property immediately simplifies (2). For an empirical measure $C_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{U_{i}}$ from a sample, the second expectation vanishes, leaving
@@
Proposition 5.1 (Witness Function Representation)
@@
Proof. Let $\mathcal{H}$ be the scalar RKHS with kernel $k$, so $\mathcal{H}^{d}$ is the $d$-fold product with $\langle g,h\rangle_{\mathcal{H}^{d}}=\sum_{j=1}^{d}\langle g_{j},h_{j}\rangle_{\mathcal{H}}$. By reproducing and closure under partials, $\partial_{u_{j}}g_{j}(u)=\langle g_{j},\,\partial_{u_{j}}k(u,\cdot)\rangle_{\mathcal{H}}$ and $g_{j}(u)=\langle g_{j},\,k(u,\cdot)\rangle_{\mathcal{H}}$. Hence
$\mathcal{T}_{C_{0}}g(u)=\sum_{j=1}^{d}\Big\langle g_{j},\;\partial_{u_{j}}k(u,\cdot)+(\partial_{u_{j}}\log c_{0}(u))\,k(u,\cdot)\Big\rangle_{\mathcal{H}}=\big\langle g,\,\xi_{C_{0}}(u)\big\rangle_{\mathcal{H}^{d}},$
with component $\xi_{C_{0}}(u)_{j}:=\partial_{u_{j}}k(u,\cdot)+s_{j}(u)\,k(u,\cdot)\in\mathcal{H}$. Thus $\xi_{C_{0}}(u)\in\mathcal{H}^{d}$.
This witness representation allows us to solve the supremum in (4) analytically. The absolute value is redundant for $C_{n}$ since $\mathbb{E}_{C_{0}}[\mathcal{T}_{C_{0}}g]=0$ by Lemma 1. By Hilbert space duality, the supremum is attained when $g$ aligns with the mean witness, converting the expression into a norm; squaring yields a closed-form V-statistic with the Copula-Stein kernel.
Theorem 1 (Closed-Form CSD)
@@
Proof. By Lemma 1,
$\text{CSD}(C_{n},C_{0})=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\left|\frac{1}{n}\sum_{i=1}^{n}\mathcal{T}_{C_{0}}g(U_{i})\right|=\sup_{\|g\|\leq 1}\left|\left\langle g,\;\mu_{n}\right\rangle_{\mathcal{H}^{d}}\right|,\quad\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\xi_{C_{0}}(U_{i}).$
$\text{CSD}^{2}(C_{n},C_{0})=\|\mu_{n}\|_{\mathcal{H}^{d}}^{2}=\left\langle\frac{1}{n}\sum_{i}\xi_{C_{0}}(U_{i}),\;\frac{1}{n}\sum_{j}\xi_{C_{0}}(U_{j})\right\rangle_{\mathcal{H}^{d}}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\langle\xi_{C_{0}}(U_{i}),\xi_{C_{0}}(U_{j})\rangle_{\mathcal{H}^{d}},$
which identifies $k_{C_{0}}(u,v):=\langle\xi_{C_{0}}(u),\xi_{C_{0}}(v)\rangle_{\mathcal{H}^{d}}$.
@@ 5.2 A practical test
$\widehat{\mathrm{CSD}}_{n}^{2}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}),$
which aligns with our asymptotic theory (Sec. 6) and our implementations (Sec. 8). We obtain critical values via the wild bootstrap: draw $w_{i}$ i.i.d. Rademacher and compute
$T^{\ast}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{i}w_{j}\,k_{C_{0}}(U_{i},U_{j}).$
Repeat $B$ times (e.g., $B=1000$) and reject for $\widehat{\mathrm{CSD}}_{n}^{2}>q_{1-\alpha}\{T^{\ast}\}$. This preserves the degeneracy and yields valid level-$\alpha$ tests; see Chwialkowski et al. (2016).
Remark 5.1
@@
Proposition 5.2 (Explicit Kernel Formula)

### v2

@@ Abstract
Kernel Stein discrepancies (KSDs) are widely used for goodness-of-fit testing, but standard KSDs can be insensitive to higher-order dependence features such as tail dependence. We introduce the Copula-Stein Discrepancy (CSD), which defines a Stein operator directly on the copula density to target dependence geometry rather than the joint score. For Archimedean copulas, CSD admits a closed-form Stein kernel derived from the scalar generator. We prove that CSD metrizes weak convergence of copula distributions, admits an empirical estimator with minimax-optimal rate $O_{P}(n^{-1/2})$, and is sensitive to differences in tail dependence coefficients. We further extend the framework beyond Archimedean families to general copulas, including elliptical and vine constructions. Computationally, exact CSD kernel evaluation is linear in dimension, and a random-feature approximation reduces the quadratic $O(n^{2})$ sample scaling to near-linear $\tilde{O}(n)$; experiments show near-nominal Type I error, increasing power, and rapid concentration of the approximation toward the exact $\widehat{\mathrm{CSD}}_{n}^{2}$ as the number of features grows.
Machine Learning, ICML
1 Introduction
@@
  5. 5.
    We provide numerical experiments demonstrating near-nominal Type I error and increasing power for the CSD test in a bivariate setting, and validating the accuracy of the random-feature approximation against the exact $\widehat{\mathrm{CSD}}_{n}^{2}$.
The remainder of the paper is structured as follows. Section 2 discusses related work. Section 3 provides background on KSDs and copulas. Section 4 introduces all the notations and assumptions used throughout the paper. Section 5 formally introduces the CSD, its operator, and kernel. Section 6 establishes its key theoretical properties. Section B extends the framework to general copulas, and Section C details its computational properties and scalable algorithms. Finally, Section 7 reports numerical experiments illustrating finite-sample testing performance and the accuracy of the random-feature approximation.
2 Related Work
@@ 2.2 Goodness-of-Fit Testing for Copulas
CSD contributes a new class of GoF tests to this literature with several key advantages. As a discrepancy metric, its value provides a quantitative measure of misspecification, whereas many classical tests only yield a p-value. Most importantly, CSD is specifically designed to be sensitive to features like tail dependence, a known challenge highlighted in financial applications (Malevergne and Sornette, 2003). While specialized tests for extreme-value dependence exist (Kojadinovic and Yan, 2011; Genest and Segers, 2009), CSD provides a general framework with theoretical guarantees on its sensitivity to tail coefficients for any copula family.
@@ 2.3 Copula Modeling and Diagnostics
3 Background: Copulas and Stein’s Method
@@ 3.2 Kernel Stein Discrepancy
4 Standing Assumptions and Notation
This section collects the standing regularity conditions used throughout the paper and fixes notation. Unless stated otherwise, all theoretical results are proved under Assumption 4.2 (A1–A5).
Distributions and Samples. A probability distribution is $P$ with density $p$. The target (null) copula distribution is $C_{0}$ with density $c_{0}$; a generic copula is $C$ with density $c$. The dd–dimensional unit hypercube is $0,1]^{d}$. An i.i.d. sample of size $n$ from a copula is $\{U_{i}\}_{i=1}^{n}$, $U_{i}\in[0,1]^{d}$. The empirical measure is $C_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{U_{i}}$. Expectations are $\mathbb{E}[\cdot]$.
Copula Framework. A copula is $C(u)$ with $u=(u_{1},\dots,u_{d})\in[0,1]^{d}$. For Archimedean copulas, $\varphi:(0,1]\to[0,\infty)$ is the generator, $\psi=\varphi^{-1}$, and $\varphi^{(k)}$ denotes the $k$th derivative. Parametric dependence parameters are denoted $\theta$. The copula score is $s(u)=\nabla\log c(u)$. Lower/upper tail coefficients are $\lambda_{L},\lambda_{U}$. For Gaussian (elliptical) copulas, $\Sigma$ is the correlation matrix and $\Phi^{-1}$ the standard normal quantile.
RKHS/Stein Framework. The base kernel is $k(u,v)$ on $0,1]^{d}\times[0,1]^{d}$ with scalar RKHS $\mathcal{H}$ and vector RKHS $\mathcal{H}^{d}$. A vector test function is $g(u)\in\mathcal{H}^{d}$. The Copula-Stein operator at target $C_{0}$ is
$\mathcal{T}_{c_{0}}g(u)=\sum_{j=1}^{d}\!\Big[\partial_{u_{j}}g_{j}(u)+g_{j}(u)\,\partial_{u_{j}}\log c_{0}(u)\Big].$ (2)
@@ 4 Standing Assumptions and Notation
Unless noted, we work with test functions generated by a boundary-conditioned RKHS (see Proposition 4.2 below), so that each component has zero trace on the coordinate faces. This eliminates boundary terms in the Stein identity on $(0,1)^{d}$.
Definition 4.1 (Copula Stein discrepancy).
@@
$\mathrm{CSD}(Q,C_{0})\;\equiv\;\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\,\mathbb{E}_{U\sim Q}\big[\,\mathcal{T}_{c_{0}}g(U)\,\big].$ (4)
Under Assumption 4.2 (A1–A4), we have $\mathrm{CSD}(C_{0},C_{0})=0$ by Lemma 5.2.
@@
Assumption 4.2 (Regularity Conditions).
@@ 3. A3:
    (Stein-kernel universality) The Copula–Stein kernel $k_{C_{0}}$ defined in Proposition 5.2 admits a bounded, continuous extension to the compact domain $0,1]^{d}$ and is $c$-universal on $0,1]^{d}$ (equivalently, its RKHS is dense in $C([0,1]^{d})$ under the sup norm).
@@ 4. A4:
    (Boundary control, concrete sufficient condition) The Stein identity for $\mathcal{T}_{C_{0}}$ holds without boundary terms. A sufficient condition (and the one we adopt for all results below) is obtained by working with a boundary-conditioned kernel: choose a base positive definite kernel $\bar{k}$ on $0,1]^{d}$ with $\bar{k}\in C^{2}$, define
    $b(u):=\prod_{j=1}^{d}u_{j}(1-u_{j}),$ (5)
    $k(u,v):=b(u)\,b(v)\,\bar{k}(u,v),$ (6)
    and let $\mathcal{H}$ be the RKHS of $k$. Then every $f\in\mathcal{H}$ vanishes on $\partial[0,1]^{d}$ and hence every $g\in\mathcal{H}^{d}$ has zero trace on each coordinate face (see Proposition 4.2). In particular, integration by parts yields the Stein identity.
@@ 5. A5:
Proposition 4.3 (Boundary-conditioned kernels yield the required trace condition).
Let $\bar{k}$ be a positive definite kernel on $0,1]^{d}$ with RKHS $\bar{\mathcal{H}}$. Let $b:[0,1]^{d}\to\mathbb{R}$ be continuous and define
$k(u,v):=b(u)\,b(v)\,\bar{k}(u,v),\qquad u,v\in[0,1]^{d}.$ (7)
Then $k$ is positive definite. Moreover, the mapping
$M_{b}:\bar{\mathcal{H}}\to\mathcal{F}([0,1]^{d}),\qquad(M_{b}\bar{f})(u):=b(u)\bar{f}(u),$ (8)
has range contained in the RKHS $\mathcal{H}$ of $k$, and every $f\in\mathcal{H}$ admits a representative that vanishes wherever $b$ vanishes. In particular, if $b(u)=0$ for all $u\in\partial[0,1]^{d}$ (e.g., $b(u)=\prod_{j=1}^{d}u_{j}(1-u_{j})$), then every $f\in\mathcal{H}$ has zero trace on $\partial[0,1]^{d}$, hence every $g\in\mathcal{H}^{d}$ has zero trace on each coordinate face.
If additionally $\bar{k}\in C^{2}([0,1]^{d}\times[0,1]^{d})$, then $k\in C^{2}([0,1]^{d}\times[0,1]^{d})$ and the reproducing identities for first-order partial derivatives used in Proposition 5.1 remain valid (on compact subsets of $(0,1)^{d}$).
Remark 4.4 (Random-feature / Nyström scaling becomes explicit).
If $\bar{k}$ admits a feature representation $\bar{k}(u,v)\approx\phi(u)^{\top}\phi(v)$, then the boundary-conditioned kernel satisfies
$k(u,v)\approx(b(u)\phi(u))^{\top}(b(v)\phi(v)).$ (9)
Thus any random-feature or Nyström approximation built for $\bar{k}$ immediately yields an approximation for $k$ by multiplying features by $b(\cdot)$. This makes the near-linear-time approximation claim constructive once $\bar{k}$ is chosen.
Strengthening used only for metrization. When needed for metrization, we additionally assume:
Assumption 4.5 (Boundary rates for metrization).
@@
Estimator (fixed): We use the V-statistic form throughout:
$\widehat{\mathrm{CSD}}_{n}^{2}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}),\quad U_{i}\in[0,1]^{d}.$ (10)
Archimedean score sum: Let $\varphi:[0,1]\to\mathbb{R}$ be the generator/score map; we write
$t(u)\;=\;\sum_{k=1}^{d}\varphi(u_{k})\,.$ (11)
Random features (Section C): We denote by $m\in\mathbb{N}$ the number of random features and by $\phi:\,[0,1]^{d}\to\mathbb{R}^{m}$ the random feature map used to approximate $k_{C_{0}}$. Given a draw $\omega\sim\Pi$ (feature distribution), write $\phi(u)=\phi(u;\omega_{1:m})$ for the $m$-dimensional feature vector.
We use a random-feature witness defined in Section C.
5 The Copula-Stein Discrepancy
The central goal of this section is to construct a dependence-aware goodness-of-fit metric relative to a target copula $C_{0}$. Throughout, we work under Assumptions 4.2 (and, when invoked for metrization results, Assumption 4.5).
@@ 5.1 From IPM to a Computable Kernel
Following the framework of integral probability metrics (IPMs) (M”uller, 1997; Zolotarev, 1983), we define a discrepancy as the worst-case difference in expectation over a class of test functions $g\in\mathcal{H}^{d}$:
$\text{CSD}(C,C_{0})$ (12)
$=\sup_{\|g\|_{\mathcal{H}^{d}}\leq 1}\left|\mathbb{E}_{U\sim C}[\mathcal{T}_{c_{0}}g(U)]-\mathbb{E}_{U\sim C_{0}}[\mathcal{T}_{c_{0}}g(U)]\right|.$ (13)
@@
Definition 5.1 (Copula-Stein Operator).
@@
Lemma 5.2 (Stein identity and boundary conditions).
@@
$\mathbb{E}_{U\sim C_{0}}\!\big[\mathcal{T}_{c_{0}}g(U)\big]=0,\quad\mathcal{T}_{c_{0}}g:=\sum_{j=1}^{d}\big(\partial_{u_{j}}g_{j}+s_{j}g_{j}\big),$ (15)
$s_{j}=\partial_{u_{j}}\log c_{0}.$ (16)
Proof is given in Appendix A.1.
Using Lemma 5.2, the zero-mean property immediately simplifies (13). For an empirical measure $C_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{U_{i}}$ from a sample, the second expectation vanishes, leaving
@@
Proposition 5.3 (Witness Function Representation).
@@
Proof is given in Appendix A.2.
This witness representation allows us to solve the supremum in (17) analytically. The absolute value is redundant for $C_{n}$ since $\mathbb{E}_{C_{0}}[\mathcal{T}_{c_{0}}g]=0$ by Lemma 5.2. By Hilbert space duality, the supremum is attained when $g$ aligns with the mean witness, converting the expression into a norm; squaring yields a closed-form V-statistic with the Copula-Stein kernel.
Theorem 5.4 (Closed-Form CSD).
@@
We have given the proof in Appendix A.3.
@@ 5.2 A practical test
$\widehat{\mathrm{CSD}}_{n}^{2}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}(U_{i},U_{j}),$ (20)
which aligns with our asymptotic theory (Sec. 6) and our implementations (Sec. C). We obtain critical values via a parametric bootstrap under the null copula $C_{0}$: for $b=1,\dots,B$, draw an i.i.d. bootstrap sample $U^{\ast(b)}_{1:n}\sim C_{0}$ (with the null parameter fixed at $\theta_{0}$, or using a plug-in estimate $\hat{\theta}$ when applicable), and compute
$T^{\ast(b)}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{C_{0}}\!\bigl(U^{\ast(b)}_{i},U^{\ast(b)}_{j}\bigr)$. (21)
Repeat $B$ times (e.g., $B=1000$) and reject for $\widehat{\mathrm{CSD}}_{n}^{2}>q_{1-\alpha}\{T^{\ast}\}$, where $q_{1-\alpha}\{T^{\ast}\}$ denotes the empirical $(1-\alpha)$-quantile of $\{T^{\ast(b)}\}_{b=1}^{B}$.
Remark 5.5.
@@
Proposition 5.6 (Explicit Kernel Formula).